# Demystifying Behavior-Based Malware Detection at Endpoints

### v1

@@ 3. Endpoint Malware Detection
Terminology. A sample is an executable program, identified by its unique SHA-256 hash. A trace is a sequence of behavioral actions (such as file accesses or process creations) performed by a sample when it is executed in a computing environment. Our work focuses on Windows environments and executable samples. We refer to an in-the-wild endpoint environment as a host. In our endpoint dataset, the anti-malware system in each host recorded the traces of samples executed by the host that were not determined to be benign or malicious in an earlier stage of the detection pipeline. Each sample can have multiple traces recorded at multiple hosts at different timestamps. A sandbox (SB) is a synthetic environment that provides tools for analyzing samples and recording their traces through controlled execution. Our sandbox dataset contains traces collected from two sandboxes from two vendors, namely Tencent HABO (habo, ) and an in-house version of Cuckoo Sandbox (cuckoo, ). In the rest of the paper, we will refer to these sandboxes as SB1 and SB2, respectively.
@@
Notation. We denote the set of samples in our dataset as $\mathbf{P}$. $P_{i}$ is the $i$-th program and $y_{i}$ is its ground truth label, where $y\!=\!0$ and $y\!=\!1$ indicate a benign and malware sample, respectively. If $P_{i}$ is a malware sample, it is also tagged with a family label $s_{i}$ that is useful for grouping samples with similar characteristics. A trace of $P_{i}$ is $x_{i}^{j}$, where $j$ denotes the execution environment, specifically, $j\!=\!0$ refers to the endpoint hosts and $j\!\in\!\{1,2\}$ refers to the sandboxes. As there are multiple endpoint traces per sample, we refer to the $k$-th one as $x_{i,k}^{0}$, and the endpoint traces of a sample are enumerated by their timestamps. The timestamp of its earliest observed trace—$x_{i,0}^{0}$—marks the first time $P_{i}$ was first seen in the wild, and $\mathbf{x}^{0}_{i,(t<h)}$ is the set of all endpoint traces recorded within $h$ hours of this.
@@
An ML-based detection model takes a trace $x_{i}$ of $P_{i}$ as its input and aims to infer $y_{i}$. We split a model $f$ into two parts: an encoder enc and a classifier $g$. The encoder produces a vector embedding $z_{i}=\text{enc}(x_{i})$ of the input trace and $f(x_{i})=g(z_{i})$ is the model’s predicted probability (score) that $P_{i}$ is malware. The predicted label $\hat{y_{i}}$ is $\hat{y_{i}}\!=\!1$ if $f(x_{i})\!\geq\!\tau$, or $\hat{y_{i}}\!=\!0$, otherwise. Here, $\tau$ is a threshold tuned based on the requirements, where a higher $\tau$ results in fewer false positives in exchange for fewer true positives. In Section 3.3, we discuss the tuning of $\tau$ in more detail.
Figure 1. An overview of endpoint malware detection.
3.1. Malware Detection Workflow
Figure 1 provides an overview of the workflow for malware detection at endpoints. We focus on behavior-based detectors deployed in step 3. There are two main phases from the vendor’s perspective: training and testing. In the training phase, the vendor collects a set of samples. These samples are often collected from large-scale repositories, either public, such as MalwareBazaar (bazaar, ), or private (ugarte2019close, ). Most prior work in this domain has followed this practice (jindal2019neurlux, ; miller2016reviewer, ; karbab2019maldy, ; dambra2023decoding, ; kuchler2021does, ). The vendor can rely on a public platform, such as VirusTotal (virustotal, ) to label these samples, or for more recent or advanced malware, can allocate malware analysts for manual labeling (yong2021inside, ). Most commonly, the collected samples are executed in sandboxes that can conveniently produce large-scale datasets of traces for training. In contrast, collecting a large-scale dataset of endpoint traces from hosts requires a telemetry infrastructure, a permissive user agreement, and a large enough user base—conditions that commercial anti-malware vendors can meet but researchers typically cannot. The training phase concludes with training an ML model on the collected traces for detecting malware from runtime behaviors. During the test phase, the model is deployed locally at the hosts. Although we view them as sequential, these phases can be interwoven as vendors continuously update their models on new data (chen2023continuous, ).
@@ 3.1. Malware Detection Workflow
In practice, vendors deploy a chain of techniques (avcomparisons, ; vmray, ) to detect if a sample $P$ encountered by a host (e.g., from an e-mail attachment) is malware—step 2 in Figure 1. First the sample is analyzed with static techniques, which operate without executing the sample. This is sufficient for classifying most existing samples (dambra2023decoding, ) either as malware ($P$ is prevented from executing) or benign ($P$ is allowed to execute). If $P$ remains undetected after step 2, it will be executed if the host demands it, e.g., by clicking on the attachment. Critically, our endpoint dataset contains only the traces recorded in the wild after this step, i.e., the traces of undetected samples. The challenges in performing behavior-based detection on such samples are unknown, as prior work has not considered it a criterion for constructing evaluation sets. If $P$ is malware, executing it leads to a real-world compromise—an infection. An endpoint detector is the last line of defense that operates at the host in real time to identify whether $P$ is malware by consuming the trace resulting from this execution—step 3. If $P$ is detected as malware during this step, anti-malware products often apply remedies such as quarantining, deleting files, or alerting (quarantine, ).
@@
Endpoint detectors must classify the sample based on an endpoint trace observed on the end host. It is not feasible to first execute the sample in a sandbox and then classify it based on the sandbox trace, as running in a sandbox (whether in the cloud (cloudsandbox, ) or on-premise (onpremsandbox, )) introduces unacceptable delays. Prior research has only evaluated ML detectors on sandbox traces, but in deployment, the detectors will be applied to endpoint traces, so their actual performance has not been evaluated in the research literature to date. We tackle this issue in this paper.
3.2. Model Training and Evaluation Scenarios
@@
Table 2 summarizes the detection scenarios, each defined along four dimensions, in our focus. The Samples dimensions indicate the step of the workflow training or testing samples originate from: 1—a broad distribution of all programs hosts encounter—or 3—a distribution of programs that remained undetected after 2. The Env. (short for environment) dimensions encode where the traces are collected—from a sandbox or endpoint hosts.
@@
Table 2. Scenarios for behavior-based malware detection.
@@
SB->SB is the scenario most prior work in our domain considers: both training and testing samples are collected from sample repositories and corpora, which represent the distribution of all samples seen in the wild, and these samples are executed in a sandbox (karbab2019maldy, ; jindal2019neurlux, ; miller2016reviewer, ; kuchler2021does, ). It is already known (and also corroborated by our results) that detectors excel (90%+ TPR) in this scenario (dambra2023decoding, ).
@@
In SB->EP and EP->EP scenarios, detectors are deployed at 3 in Figure 1, i.e., on the endpoint traces of undetected samples. In SB->EP, the detector is trained only on sandbox-based data (same as SB->SB) but deployed for endpoint detection. In EP->EP, the model is trained on the real-world endpoint traces of undetected samples. This departs from the norm in published research, as most prior detectors have relied on sandbox traces for training. We study this scenario to assess the success of a model explicitly trained for endpoint detection. These scenarios represent a trade-off between practicality (SB->EP is easiest to implement) and expected detection performance (where EP->EP is favored).
3.3. Success Metrics for Endpoint Detection
@@
A conventional detector aims to detect as many malware samples as possible, i.e., to maximize the true-positive rate (TPR), while minimizing the number of benign samples misclassified, i.e., the false-positive rate (FPR). The defender can control the TPR and FPR by tuning the detection threshold $\tau$; a higher $\tau$ implies a lower TPR and FPR. Following prior work (arp2014drebin, ), we evaluate our models regarding their TPR @1% FPR on the test sets (reported as TPR). Additionally, we report the area under the TPR-FPR curve, i.e., the receiver operating characteristic curve, to gauge a model’s overall ability to separate malware and benign classes (reported as AUC). Note that a uniform random predictor, regardless of the class proportions in the test set, would achieve 1% TPR @1% FPR and 50% AUC.
@@
Measuring TPR and FPR in SB->SB is straightforward as there is a one-to-one mapping from samples to traces. However, in endpoint detection (SB->EP and EP->EP), there is a one-to-many mapping as multiple hosts may execute the same sample at different times, and these resulting traces often have high variability (avllazagaj2021malware, ). For example, a few malware samples have over 400 traces, each corresponding to a real-world infection. We present a histogram of the number of traces per sample in Figure 6(a). To obtain a one-to-one mapping, we consider “an average trace” by selecting one random endpoint trace of each sample and measuring the TPR (averaged over 100 runs). This gives us a fair way to compare performance across different scenarios. We refine this metric further based on practical considerations. 67% of traces of an average malware sample (52% for benign) are seen within 24 hours of its first execution at a host. As this ratio drops rapidly after the first day (see Figure 6(b)) and malware samples die quickly within days (yong2021inside, ), we only randomly select from the first-day traces—$\mathbf{x}^{0}_{i,(t<24)}$—unless stated otherwise.
4. Technical Setup
4.1. Datasets
@@
Table 3. A summary of the samples in our datasets.
@@
Endpoint (EP) Dataset. Provided by Avllazagaj et al. (avllazagaj2021malware, ), we use a dataset of program traces recorded on real Windows hosts of a commercial anti-malware vendor from over 100 countries between January and July 2018. The vendor did not know whether the samples were benign or malicious at the time of execution. The samples were executed by the users, who interacted with them naturally, and the vendor’s behavior-based component recorded the traces in a last-ditch effort to discover unknown threats. We have reprocessed this dataset specifically for our problem and relabeled it by querying VirusTotal (virustotal, ) for more accurate labels. Our processed dataset contains around 1M execution traces from 900 malware samples and 25K benign samples. Each trace includes high-level actions (file, registry, process, and mutex actions) of a sample that is executed until its termination. To our knowledge, this is the only dataset that allows us to evaluate endpoint detection at scale.
@@
Sandbox (SB) Dataset. Studying SB->EP realistically requires collecting a sandbox dataset contemporaneous to our EP dataset. Because the sandbox dataset will serve as the training set and the EP dataset as the test set, we must collect samples seen in or before 2018 to respect causality (pendlebury2019tesseract, ). Moreover, we cannot detonate samples in a sandbox, as it is commonly done (karbab2019maldy, ; jindal2019neurlux, ; miller2016reviewer, ; kuchler2021does, ), as behaviors of old programs today would differ from their original behaviors. This is because malware stops functioning when, for example, its remote infrastructure dies (yong2021inside, ) or it starts behaving differently over time (avllazagaj2021malware, ). Our solution is collecting traces from VirusTotal (virustotal, ), where third-party sandbox vendors publicly share behavior reports on samples. To this end, we curate a list of SHA-256 hashes of Windows samples from popular public malware detection corpora, including EMBER (anderson2018ember, ), and SOREL (harang2020sorel20m, ), released in 2017 or 2018. We then query VirusTotal with these hashes to collect their sandbox traces when available. We discard traces that came more than six months after the sample was first seen to capture close-to-original behaviors. This process results in traces from two sandbox vendors for around 110K malware and 40K benign samples. These vendors are well-known in their countries of origin (SB1 is from China, and SB2 is from the USA). There is a class imbalance as vendors on VirusTotal are more inclined to share traces of malware samples.
Rank Test Set Trained on SB1 Trained on SB2 Trained on EP SB1->SB1 SB1->SB2 SB1->EP SB2->SB1 SB2->SB2 SB2->EP EP->EP NGR HYB ATT NGR HYB ATT NGR HYB ATT NGR HYB ATT NGR HYB ATT NGR HYB ATT NGR HYB ATT SB1 95.0 94.2 93.2 60.7 53.7 56.7 11.2 8.0 4.5 48.5 31.8 32.7 85.5 59.2 83.7 14.2 7.2 12.3 43.5 34.0 25.7 SB2 94.0 93.1 92.2 70.4 61.2 64.3 13.9 7.1 5.5 17.6 10.0 14.7 92.6 91.4 90.8 7.5 6.3 11.4 43.4 37.6 26.4 EP 93.2 93.1 87.8 63.5 48.0 52.9 16.7 10.2 8.3 34.5 16.6 15.7 79.3 79.5 88.7 17.3 10.9 13.1 49.5 43.5 42.8
@@
Table 4. The performance (TPR%) of three ML approaches (NGR, HYB, ATT) in seven detection scenarios (based on Table 2). In each row, the models are ranked based on the test set, and the average TPR of top-20 models is reported (see Table 15 for AUC).
Trace Standardization. Our EP traces contain high-level information that the vendor can seamlessly collect and analyze at endpoint hosts, such as file, process, and mutex creations, as well as registry key creations and deletions. Sandboxes also record lower-level information (such as memory dumps (sandboxmemory, )) as they are less computationally constrained. Moreover, our traces from three sources (SB1, SB2, and EP) have different formats and conventions. This prompts us to implement routines to keep, in our sandbox traces, only the type of information available in EP traces and covert all traces to a standardized format (see Appendix A for details). As shown in Section 5, the information remaining in sandbox traces is still enough to obtain performance comparable to prior work ($\sim$95TPR and 99% AUC) in the SB->SB setting. This standardization is also critical for our leaderboard design (Section 8).
@@
Temporal Splitting and Labeling. We split our datasets into two portions based on the timestamps of the samples (the first-seen dates) to ensure that the train and test samples are temporally disjoint, avoiding a common pitfall in prior work (pendlebury2019tesseract, ; arp2020and, ). Samples seen before April 1st, 2018, along with their traces, are in the Training portion, and the samples seen after that date are in the Testing portion, on which we never train. As labels of samples are known to change mildly over time (labeldynamics, ), we make the best effort to assign the training samples historical labels that were available before April 1st, 2018. We could find such historical labels for $\sim$24and 100% of the samples in the EP and SB training portions, respectively. For testing samples, we query VirusTotal to obtain the most recent detection reports and label samples detected by over five anti-malware engines as malware, following the advice in (labeldynamics, ). Further, we use AVClass2 (sebastian2020avclass2, ) to assign family labels to our malware samples based on their detection reports. We tag malware samples that did not receive a family label as Generic malware. Overall, this methodology ensures that our evaluation is realistic and reflects the conditions the anti-malware vendor had to work under when our endpoint data was collected.
4.2. Machine Learning Details

### v2

@@ III Endpoint Malware Detection
Terminology. A sample is an executable program identified by its unique SHA-256 hash. A trace is a sequence of behavioral actions (such as file accesses or process creations) performed by a sample when it is executed in a computing environment. We refer to an in-the-wild endpoint environment as a host. Our work focuses on Windows hosts and executable samples. In our endpoint dataset, the anti-malware system in each host recorded the traces of samples executed by the host that were not determined to be benign or malicious at an earlier stage of the detection pipeline. Each sample can have multiple traces recorded at multiple hosts at different times. A sandbox (SB) is a synthetic environment that provides tools for analyzing samples and recording their traces with controlled executions. Our sandbox dataset contains traces collected from two commercial sandboxes: Tencent HABO [47] and VirusTotal [48], which we will refer to as SB1 and SB2, respectively.
@@
Notation. We denote an individual sample in our dataset as $P_{i}$ and its ground truth label as $y_{i}$, where $y_{i}\!=\!0$ and $y_{i}\!=\!1$ indicate a benign and malware sample, respectively. If $P_{i}$ is a malware sample, it is also tagged with a family label $s_{i}$ that is useful for grouping samples with similar characteristics. A trace of $P_{i}$ is $x_{i}^{j}$, where $j$ denotes the execution environment, specifically, $j\!=\!0$ refers to the endpoint hosts and $j\!\in\!\{1,2\}$ refers to the sandboxes. As there are multiple endpoint traces per sample, we refer to the $k$-th one as $x_{i,k}^{0}$, and the endpoint traces of a sample are enumerated by their timestamps. The timestamp of its earliest observed trace—$x_{i,0}^{0}$—marks the first time $P_{i}$ was first seen in the wild, and $\mathbf{x}^{0}_{i,(t<h)}$ is the set of all endpoint traces recorded within $h$ hours of this.
@@
An ML-based detection model takes a trace $x_{i}$ of $P_{i}$ as its input and aims to infer $y_{i}$. We split a model $f$ into two parts: an encoder enc and a classifier $g$. The encoder produces a vector embedding $z_{i}=\text{enc}(x_{i})$ of the input trace and $f(x_{i})=g(z_{i})$ is the model’s predicted probability (score) that $P_{i}$ is malware. The predicted label $\hat{y_{i}}$ is $\hat{y_{i}}\!=\!1$ if $f(x_{i})\!\geq\!\tau$, or $\hat{y_{i}}\!=\!0$, otherwise. Here, $\tau$ is a tunable threshold, where a higher $\tau$ reduces false positives in exchange for fewer true positives. In Section III-C, we discuss the tuning of $\tau$ in more detail.
Figure 1: An overview of endpoint malware detection.
III-A Malware Detection Pipeline
@@
Figure 1 provides an overview of the pipeline for malware detection at endpoints. In practice, vendors deploy a chain of techniques [3, 9] to detect if a sample $P$ encountered by a host (e.g., from an e-mail attachment) is malware—step 2 and 3. If one technique in the pipeline fails to classify the sample as malware or benign conclusively, the next one is invoked. First, it deploys high-precision methods (blocklists and signatures), then less precise static ML approaches [49, 4], and finally, dynamic and behavioral analysis. We focus on ML-based behavioral malware detectors deployed in step 3. Static techniques operate without executing the sample and are deployed first as they are more efficient and secure—step 2. This is sufficient for classifying most existing samples [26] either as malware ($P$ is prevented from executing) or benign ($P$ is allowed to execute). If $P$ remains undetected after this step, it will be executed if the host demands it, e.g., by clicking on the attachment. If $P$ is malware, executing it leads to a real-world compromise—an infection. An endpoint detector is the last line of defense that operates at the host in real-time to identify whether $P$ is malware by consuming the trace resulting from this execution—step 3. If $P$ is detected as malware during this step, anti-malware products often apply remedies such as quarantining, deleting files, or alerting [8].
@@
We identified two issues regarding this pipeline in past research. First, prior work constructed evaluation sets for behavioral malware detectors without considering their position in an end-to-end pipeline. As a result, the challenges in performing behavioral detection on samples that bypassed static methods are unknown. Second, endpoint detectors must classify the sample based on an endpoint trace observed at the end host. It is not feasible to first execute the sample in a sandbox and then classify it based on the sandbox trace, as running in a sandbox (whether in the cloud [50] or on-premise [14]) introduces unacceptable delays. Prior work has only evaluated ML detectors on sandbox traces, but in deployment, the detectors will be applied to endpoint traces, so their actual performance has not been accurately evaluated. We tackle these issues by performing realistic evaluations using our endpoint dataset that contains only the traces of undetected samples—which reached the step 3 in the wild.
Assumptions and Limitations. Security vendors such as VMRay [9], Palo Alto Networks [51], and Kaspersky [52] use multi-tiered solutions that chain static and dynamic analysis in their anti-malware products. We do not have access to these proprietary pipelines. Instead, we consider a streamlined one in Figure 1 based on public documentation by these vendors. Such a pipeline should ideally be evaluated end-to-end, starting with a large pool of samples from the wild, where each component sees only the samples that previous components cannot classify. We are unable to perform such an evaluation as we do not have raw binaries (to extract static features) of our endpoint samples (we could find less than 10% of samples from public sources, mostly malware). Prior malware detection datasets (e.g., SOREL [53]) only share vectorized static features and not the binaries, preventing the simulation of behavioral components. However, we can still realistically isolate and evaluate behavioral ML methods at the pipeline’s last stage as our endpoint dataset contains samples that bypassed a real-world product’s static components. We recommend future work to collect datasets (with static and dynamic features) that allow the evaluation of the whole pipeline end-to-end or each component according to its position.
Moreover, a pipeline’s components constantly evolve in practice. For example, a sample from a new malware family might initially bypass static methods. Eventually, this family can be caught at step 2 with new static signatures or updates to static ML models. Evaluating this requires a dataset with a large temporal span, whereas our endpoint data covers only six months (insufficient for significant changes [25]). Consequently, we perform our evaluation with fixed behavioral ML models that are trained at a particular timestamp and tested on future samples that bypass static methods.
III-B Model Training and Evaluation Scenarios
There are two main phases for deploying an ML-based endpoint detector: training and testing. In the training phase, the vendor collects a set of samples. These samples are often collected from large-scale repositories, either public, such as MalwareBazaar [27], or private [54]. Most prior work in this domain has followed this practice [6, 55, 5, 26, 28]. The vendor can rely on a public platform, such as VirusTotal [48], to label these samples or employ malware analysts in more advanced cases [56]. Most commonly, the collected samples are executed in sandboxes that can conveniently produce large-scale datasets of traces for training. In contrast, collecting a large-scale dataset of endpoint traces from hosts requires a telemetry infrastructure, a permissive user agreement, and a large enough user base—conditions that commercial anti-malware vendors can meet but researchers typically cannot. The collected traces are used to train an ML model that can detect malware from runtime behaviors. During the test phase, the model is deployed locally at each host. Training and testing phases can also be interwoven as vendors continuously update their models on new data [39].
@@ III-B Model Training and Evaluation Scenarios
Table II summarizes the detection scenarios, each defined along four dimensions, considered in our work. The Samples dimensions indicate the step of the pipeline training or testing samples originate from: 1—a broad distribution of all programs hosts encounter—or 3—a distribution of programs that remained undetected after 2. The Env. (short for environment) dimensions encode where the traces are collected—from a sandbox or endpoint hosts.
@@
TABLE II: Scenarios for behavioral malware detection.
@@
SB->SB is the scenario most prior work in our domain considers: both training and testing samples are collected from sample repositories and corpora, which represent the distribution of all samples seen in the wild, and these samples are executed in a sandbox [5, 6, 55, 28]. It is already known (and also corroborated by our results) that detectors excel (90%+ TPR) in this scenario [26].
@@
In SB->EP and EP->EP scenarios, detectors are deployed at 3 in Figure 1, i.e., on the endpoint traces of undetected samples. In SB->EP, the detector is trained only on sandbox-based data (same as SB->SB) but deployed for endpoint detection. In EP->EP, the model is trained on the real-world endpoint traces of undetected samples. This departs from the norm in published research, as most prior detectors have relied on sandbox traces for training. We study this scenario to assess the success of a model trained explicitly for endpoint detection. These scenarios represent a trade-off between practicality (where SB->EP is favored) and detection performance (where EP->EP is favored).
III-C Success Metrics for Endpoint Detection
@@
A conventional detector aims to detect as many malware samples as possible, i.e., to maximize the true-positive rate (TPR), while minimizing the number of benign samples misclassified, i.e., the false-positive rate (FPR). The defender can control the TPR and FPR by tuning the detection threshold $\tau$; a higher $\tau$ implies a lower TPR and FPR. Following prior work [57], we evaluate our models regarding their TPR @1% FPR on the test sets (reported as TPR). Additionally, we report the area under the TPR-FPR curve, i.e., the receiver operating characteristic curve, to gauge a model’s overall ability to separate malware and benign classes (reported as AUC). Note that a uniform random predictor, regardless of the class proportions in the test set, would achieve 1% TPR @1% FPR and 50% AUC.
@@
Measuring TPR and FPR in SB->SB is straightforward as there is a one-to-one mapping from samples to traces. However, in endpoint detection (SB->EP and EP->EP), there is a one-to-many mapping as multiple hosts may execute the same sample at different times, and these resulting traces often have high variability [15]. For example, a few malware samples have over 400 traces, each corresponding to a real-world infection (see Figure 7(a)). To obtain a one-to-one mapping, we consider “a representative trace” by randomly selecting one endpoint trace of each sample to create an EP test set on which we measure a model’s TPR. We repeat this process 100 times and report the average TPR over these runs as the final TPR. This gives us a fair and unambiguous way to compare performance across scenarios. We refine this metric further based on practical considerations. 67% of traces of an average malware sample (52% for benign) are seen within 24 hours of its first execution at a host. As this ratio drops rapidly after the first day (see Figure 7(b)) and malware samples die within days [56], we only randomly select from the first-day traces—$\mathbf{x}^{0}_{i,(t<24)}$—unless stated otherwise.
IV Technical Setup
IV-A Datasets
@@
TABLE III: A summary of the samples in our datasets.
@@
Endpoint (EP) Dataset. Provided by Avllazagaj et al. [15], we use a dataset of program traces recorded on real Windows hosts of a commercial anti-malware vendor from over 100 countries between January and July 2018. The vendor did not know whether the samples were benign or malicious at the time of execution. The samples were executed by the users, who interacted with them naturally, and the vendor’s behavioral component recorded the traces in a last-ditch effort to discover unknown threats. We have reprocessed this dataset for our problem and relabeled it by querying VirusTotal [48] for more accurate labels. Our processed dataset contains around 1M execution traces from 900 malware and 25K benign samples. Each trace includes high-level actions (file, registry, process, and mutex actions) of a sample that is executed until its termination. To our knowledge, this is the only dataset that allows us to evaluate endpoint detection at scale.
@@
Sandbox (SB) Dataset. A realistic study of SB->EP requires a sandbox dataset contemporaneous to our EP dataset. Because the sandbox dataset will serve as the training set and the EP dataset as the test set, we must collect samples seen in or before 2018 to respect causality [38]. Moreover, we cannot detonate samples in a sandbox, as it is commonly done [5, 6, 55, 28], as behaviors of old programs today would differ from their original behaviors. This is because malware stops functioning when, for example, its remote infrastructure dies [56] or it starts behaving differently over time [15]. Our solution is collecting traces from VirusTotal [48], where third-party sandbox vendors publicly share behavior reports on samples. To this end, we curate a list of SHA-256 hashes of Windows samples from popular public malware detection corpora, including EMBER [49], and SOREL [53], released in 2017 or 2018. We then query VirusTotal with these hashes to collect their sandbox traces when available. We discard traces that came more than six months after the sample was first seen to capture close-to-original behaviors. This results in traces from two sandbox vendors for around 110K malware and 40K benign samples. These vendors are well-known in their countries of origin (SB1 is from China, and SB2 is from the USA). There is a class imbalance as vendors on VirusTotal are more inclined to share traces of malware samples.
Sel. Test Set Trained on SB1 Trained on SB2 Trained on EP SB1->SB1 SB1->SB2 SB1->EP SB2->SB1 SB2->SB2 SB2->EP EP->EP NGR HYB ATT NGR HYB ATT NGR HYB ATT NGR HYB ATT NGR HYB ATT NGR HYB ATT NGR HYB ATT SB1 95.0 94.2 93.2 60.7 53.7 56.7 11.2 8.0 4.5 48.5 31.8 32.7 85.5 59.2 83.7 14.2 7.2 12.3 43.5 34.0 25.7 SB2 94.0 93.1 92.2 70.4 61.2 64.3 13.9 7.1 5.5 17.6 10.0 14.7 92.6 91.4 90.8 7.5 6.3 11.4 43.4 37.6 26.4 EP 93.2 93.1 87.8 63.5 48.0 52.9 16.7 10.2 8.3 34.5 16.6 15.7 79.3 79.5 88.7 17.3 10.9 13.1 49.5 43.5 42.8
@@
TABLE IV: The performance (TPR%) of three ML approaches (NGR, HYB, ATT) in seven detection scenarios (based on Table II). The models are selected using the test set in each row, and the average TPR of the top 20 models is reported.
Trace Standardization. Our traces from three sources (SB1, SB2, and EP) have different formats, contents, and conventions. We first pre-process our sandbox traces to keep only the type of behaviors recorded in our EP traces, such as file, process, mutex, and registry key creations. We then convert all traces to a single standardized format (see Section A-A for details). As shown in Section V, the resulting sandbox traces allow us to train models with comparable performance to prior work ($\sim$95TPR and 99% AUC) in the SB->SB scenario.
@@
Temporal Splitting and Labeling. We split our datasets into two portions based on the timestamps of the samples (the first-seen dates) to ensure that the train and test samples are temporally disjoint, avoiding a common pitfall in prior work [38, 42]. Samples seen before April 1st, 2018, and their traces are in the Training portion, and the samples seen after that are in the Testing portion, on which we never train. As labels of samples are known to change mildly over time [58], we make the best effort to assign the training samples historical labels that were available before April 1st, 2018. We could find such historical labels for $\sim$24and 100% of the samples in the EP and SB training portions, respectively. For testing samples, we query VirusTotal to obtain the most recent detection reports and label samples detected by over five anti-malware engines as malware, following the advice in [58]. Further, we use AVClass2 [59] to assign family labels to our malware samples based on their detection reports. We tag malware samples that did not receive a family label as Generic malware. This methodology ensures that our evaluation is realistic and reflects the conditions the anti-malware vendor had to work under when our endpoint data was collected.
IV-B Machine Learning Details